OpenStudy (anonymous):

lim(x to 0)[(cosx)/x]=

myininaya (myininaya):

its undefined from the right it goes to infinity from the left it goes to -infinity

OpenStudy (anonymous):

Use the squeeze theorem \[\lim_{x\to\infty}\cos((\pi/x)-x)/x = \lim_{x\to\infty}sin(x)/x = 0 \] It follows that; \[\lim_{x\to\infty}\cos(x)/x = 0 \]

OpenStudy (anonymous):

Sorry bout that It is undefined, myininaya is right

OpenStudy (anonymous):

the limit doesnt exist

OpenStudy (anonymous):

yep it doesn't exist would be an accurate way of saying it.

OpenStudy (anonymous):

Be careful. As x approaches zero, (cos x)/x gets larger and larger. So limit is +infinity.